If 150 is increased by 175 percent, what is the resulting number? Plz help
Found 2 solutions by jhunjiro, ikleyn:
Answer by jhunjiro(67)

Converting percent to real number.
just divide the percent to 100
example: 150 percent = 1.50
so 175 percent is equal to 1.75
to get how it increase.. you just need to multiply
150*1.75
The Answer is 262.5

Answer by ikleyn(52788)

.
If 150 is increased by 175 percent, what is the resulting number? Plz help
~~~~~~~~~~~~~~~~~~~~

The starting number is 150.


The increment is  1.75*150 = 262.5.


If 150 is increased by 175 percent, the resulting number will be

   the starting number + the increment = 150 + 262.5 = 412.5.    ANSWER

Solved.

The solution by the other tutor is incorrect.